OpenStudy (anonymous):

Which is the center of the circle whose equation is (x - 3)2 + (y - 4)2 = 121? (−3, 4) (3, −4) (−4, 3) (3, 4)

OpenStudy (anonymous):

3, 4

OpenStudy (anonymous):

thanks, how did you get that ?

OpenStudy (anonymous):

(x – h)^2 + (y – k)^2 = r^2

OpenStudy (anonymous):

(x-a)^2+(y-b)^2= r^2 is the standard form of a circle with a,b as the centre and r as the radius!

OpenStudy (anonymous):

use general formula (x-a)^2 + (x-b)^2 = r^2 whre centre = (-a,-b) just compare this with your expression and its easy

OpenStudy (anonymous):

sorry! = (a,b) is the centre